# The Impact of Large Language Models (LLMs) on Code Review Process

### v1

@@ Abstract
[Context:] Large language models (LLMs) have recently gained prominence in the field of software development, significantly boosting productivity and simplifying teamwork. Although prior studies have examined task-specific applications, the phase-specific effects of LLM assistance on the efficiency of code review processes remain underexplored. [Objectives:] This research investigates the effect of GPT on GitHub pull request (PR) workflows, with a focus on reducing resolution time, optimizing phase-specific performance, and assisting developers. [Methods:] We curated a dataset of 25,473 PRs from 9,254 GitHub projects and identified GPT-assisted PRs using a semi-automated heuristic approach that combines keyword-based detection, regular expression filtering, and manual verification until achieving 95% labeling accuracy. We then applied statistical modeling, including multiple linear regression and Mann–Whitney U test, to evaluate differences between GPT-assisted and non-assisted PRs, both at the overall resolution level and across distinct review phases. [Results:] Our research has revealed that early adoption of GPT can substantially boost the effectiveness of the PR process, leading to considerable time savings at various stages. Our findings suggest that GPT-assisted PRs reduced median resolution time by more than 60% (9 hours compared to 23 hours for non-assisted PRs). We discovered that utilizing GPT can reduce the review time by 33%, and the waiting time before acceptance by 87%. Analyzing a sample dataset of 300 GPT-assisted PRs, we discovered that developers predominantly use GPT for code optimization (60%), bug fixing (26%), and documentation updates (12%). [Conclusion:] Our findings highlight the potential of GPT to accelerate code reviews, improve code quality, and address persistent bottlenecks. This research sheds light on the impact of the GPT model on the code review process, offering actionable insights for software teams seeking to enhance workflows and promote seamless collaboration.
@@ 1 Introduction
Modern software development is heavily based on collaborative workflows, with Pull Requests (PR) serving as a central mechanism to integrate code changes [1, 2, 3]. A PR goes through multiple phases, including submission, review, change, and merging, each requiring coordination between contributors and reviewers [1]. Effective PR management is crucial to maintaining software quality, but it remains a bottleneck in many projects due to long review times, inconsistent feedback, and inefficient communication [4]. Developers often experience delays due to unclear review comments, excessive revisions, and prolonged waiting times before changes are accepted [5].
Inefficiencies in the software development cycle, such as long review times, can lead to frustration among team members, decreased productivity, and even project delays [6]. As codebases become increasingly intricate, there is a growing need for sophisticated automation tools to facilitate the code review process and improve collaboration and teamwork [7]. In parallel, LLMs, such as Generative Pre-trained Transformers (GPTs) [8] have rapidly evolved, demonstrating strong capabilities in code generation, bug detection, and documentation assistance [9, 10]. These models are capable of suggesting enhancements to existing code, resolving unclear logical structures, and automating repetitive tasks, thereby becoming indispensable tools for software engineers.
Despite the increasing adoption of LLM-powered tools in development workflows, a systematic investigation into their impact on PR processes is lacking. Previous research on the use of LLM in the PR context [9, 11, 12, 13] has focused mainly on task-specific applications, such as commit message generation [14] and documentation support [15], while neglecting to explore their phase-specific impact across the entire PR lifecycle. Furthermore, while LLM-generated suggestions can be beneficial, it is necessary to quantify their effectiveness in improving review speed and reducing PR merge time.
@@
Toward this goal, we examine the PRs that software developers actively engaged with GPT to seek recommendations for code review tasks or problem-solving. We refer to these as GPT-assisted PRs throughout the paper. In contrast, GPT non-assisted PRs are those where developers either do not use GPT at all or use it strictly for implementation tasks (i.e., making an API call to a GPT model), without relying on it for code review insights or decision making. Our research investigates how GPT-assisted PRs influence the duration and quality of the code review process. For this purpose, we start by curating a dataset of 25,473 PRs from 9,254 projects on GitHub. Next, we extract six features (project_age, no_commits, files_added, files_deleted, files_changed, time_to_merge) to characterize projects, PRs, and code reviews. We then analyze how these factors influence the review and resolution times [16]. Finally, we also investigate which development tasks involve the use of GPT-assisted PRs throughout the PR phases, aiming to identify where such assistance is most effectively integrated. In summary, our aim is to answer the following three research questions:
@@ 2 Related Work
Challenges in Modern Code Reviews (MCR) and Collaboration. Code reviews are a critical aspect of modern software development, ensuring code quality, maintainability, and adherence to best practices. However, the process comes with several challenges that can affect efficiency, collaboration, and overall development workflow. For example, long waiting times where PRs often remain unreviewed for extended periods, slowing down development. Kudrjavets et al. [5] reported that between 29% and 63% of the overall code review lifetime is spent waiting for the accepted changes to merge. Bacchelli and Bird [3] reveal that scheduling and time issues also appeared challenging; yet, understanding the code and comments takes most of the review time. According to these authors, code reviews tend to be less focused on detecting defects than expected and, instead, provide additional advantages, such as knowledge transfer, increased team awareness, and creation of alternative solutions to problems. Davila and Nunes [17] also reported that reviewers have found it difficult to understand the code, which prompts the need for enhancements in visualizing code changes to facilitate understanding. However, a recent study has shown that ChatGPT is an alternative to generating some kind of explanations during MCR [18]. Davila et al. [15] identified five challenges to adopting generative AI in MCR: trustworthiness, lack of context, misleading review comments, security and privacy, and token limits. Although prior studies have explored aspects of the use of GPT in software development, they do not provide a systematic analysis of its impact on waiting times and efficiency in all phases of the PR workflow. Our study addresses this gap by focusing on its influence during specific code review phases.
@@
Factors Affecting PR Lifetime and Latency. Studying the factors that influence PR latency is crucial for improving software development efficiency, collaboration, and overall code quality. Understanding what causes delays in PR review and merging can lead to optimizations that benefit both developers and organizations [4], [19], [16], [20], and [21]. Zhang et al. [4] identified that the number of commits at open time is negatively correlated with PR merging but shifts to a positive correlation at close time. Rose [19] built a Multiple Linear Regression Model (MLR) that indicates that the PR size has a statistically significant effect on the review time, as the positive value of the regression coefficient implies that as the PR size increases, the review time for PRs also increases. Moreira Soares et al. [16] further emphasized that PRs possess some structural characteristics, such as the number of modified files, the number of commits, and the number of changed lines of code. Moreira Soares et al. [16] concluded that the combination of structural characteristics may increase the chances of PRs having a very short and long lifetime.
@@
Although prior research explores factors affecting PR latency (e.g., size, number of commits), there is limited research into how GPT can have an effect on PR resolution time. Khatoonabadi et al. [20] developed a machine learning model to predict the latency of the initial responses of the maintainers and contributors after receiving the first feedback from the maintainer. These authors highlighted that the timing of submissions and feedback are factors that have a significant influence on response waiting times. Their approach, which provides estimated waiting times, helps open-source projects collaborate better by allowing maintainers and contributors to anticipate and address potential delays in PR reviews. This study complements the aforementioned work by offering empirical insights into GPT usage beyond commit messages, specifically examining detailed time reductions in key phases such as review and change.
@@
The Application of LLMs in Code Reviews. LLM-powered tools can offer services designed to streamline various code review-related development tasks. Research has been conducted on the impact of AI-driven generative tools in this context [9], [14], [15], [13], [22], [23], [24], [25]. For instance, Davila et al. [15] conducted a literature review on the application of AI-based generative tools in a code review context and highlighted that AI-based tools can positively impact code review efficiency by reducing reviewer workload and review duration and increasing the likelihood of acceptance in GitHub projects. Xiao et al. [14] observed that PRs assisted by Copilot have a 1.57 times higher likelihood of being merged compared to those not assisted by Copilot for PR. Tufano et al. [9] reported how developers use ChatGPT to automatically generate a commit message and a PR/issue description. Rasheed et al. [13] highlighted the potential of LLM-powered AI agents to considerably improve conventional code review procedures. These agents provide a twofold advantage: they enhance code quality and improve developer learning.
@@
Vijayvergiya et al. [22] shared the experience in developing, implementing, and evaluation AutoCommenter, a LLM-powered code review tool that enforces coding best practices and has a positive impact on the developer’s workflow. Yu et al. [23] proposed Carllm (Comprehensibility of Automated Code Review using LLMs) that improves the comprehensibility of automated code review. Hao et al. [24] studied the role of ChatGPT in collaborative coding by analyzing the shared conversation between developers and ChatGPT in GitHub PRs and issues, finding that the most frequent requests involve code generation, conceptual understanding, how-to guidance, issue resolution, and code review. Studies on GPT have examined particular tasks (e.g., code generation, commit message generation, or bug fixing) but lack a holistic view of its phase-specific impacts during the PR life cycle. Our research fills this void by examining the use and impact of GPT on specific phases, including the review process and the waiting period.
@@ 3 Study Setup
Figure 1: An overview of our study setup that describes a semi-automated approach for identifying, detecting, and mapping GPT-assisted PRs to the phases in which GPT is used
In this paper, we investigate the impact of using GPT as an assistant tool in the PR code review process. Specifically, we examine how GPT influences the overall PR resolution time and how its effects vary between different phases of the review workflow. We also aim to identify the types of review tasks for which developers actively seek GPT support. Figure 1 presents an overview of our study design. We begin with an automated procedure to identify potential GPT-assisted PRs, followed by a semi-automated verification step to confirm the actual use of GPT during the review. Finally, we map the verified GPT-assisted PRs to their corresponding review phases. The remainder of this section provides a detailed explanation of each step in our methodology.
@@ 3.1 Searching Potential GPT-assisted PRs
To identify relevant PRs for our analysis, we rely on GitHub, which adopts a PR-based approach to submit, review, and merge code contributions [17]. We use the GitHub REST API [26] to retrieve a broad set of candidate PRs in which the assistant of a GPT model could have been used during the review process.
@@ 3.1.1 Searching
Following a similar strategy to that of  Tufano et al. [9], we conduct keyword-based searches for mentions of ”GPT” in four main PR fields.
@@ 1. $\blacksquare$
    Project Name: Although some GPT-focused projects may not actually use GPT for PR reviews, we still record these mentions before filtering out irrelevant cases [27].
@@ 2. $\blacksquare$
    PR Title: The use of GPT may be evident in the PR title (e.g., refactored using GPT suggestions [28]).
@@ 3. $\blacksquare$
    Body+Comments: A PR description and/or comments can refer to GPT explicitly (e.g., code generated by GPT [29]).
@@ 4. $\blacksquare$
    File(s) Name: References in filenames (e.g., chatgpt_review.md [30]) can indicate GPT-generated artifacts.
@@ 3.1.2 Filtering
To avoid projects with minimal activity or that lack meaningful collaboration, we keep only projects that meet the following criteria: (1) the repository of the PR has at least 10 stars [31], to remove toys or trivial projects. (2) at least two contributors are involved in the review process, to exclude self-reviewed PRs.
3.1.3 Downloading
@@ 3.2 Identifying GPT-assisted PRs
In this semi-automated step, our goal is to define the heuristics that indicate the use of GPT to classify PRs as GPT-assisted or GPT-non-assisted. By doing this, we exclude false positives (i.e., PRs that mentioned GPT as discussed in Section 3.1 but did not use the GPT model to assist in the code review process). This step ensures that the PRs used for our analysis are PRs where GPT models are used in the code review process, rather than just being part of the project code.
3.2.1 Random Samples
With the goal of discovering patterns related to the use of GPT in PRs, we randomly sampled a total of 450 GitHub PRs to examine indicators of GPT assistance in code review activities. Three authors independently inspected 150 unique PRs in three separate rounds, analyzing project names, PR titles, modified files, and PR bodies. In Round 1, we reviewed the first set of 150 PRs and cross-validated our observations to ensure that the identified patterns consistently reflected GPT-assisted contributions. In Rounds 2 and 3, we applied the same procedure to two additional sets of 150 PRs each. After each round, the reviewers met to consolidate recurring patterns, clarify ambiguous cases, and resolve classification disagreements through discussion until consensus was reached. The findings of Round 1 are presented in Table 1. In particular, most PRs were ultimately classified as GPT-non-assisted, as GPT was often integrated into the software (e.g., through GPT APIs) but not explicitly used to support the code review itself.
@@ 3.2.2 Defining Heuristics
We defined four heuristics to analyze and label the presence of GPT-assisted PRs, as outlined in Table 2. These heuristics are reinforced by a regular expression-based technique (detailed in the appendix), which automatically classifies the collected PR instances as either GPT-assisted or GPT-non-assisted based on the identified patterns: H1: Project name, H2: PR title, H3: Modified files, H4: PR body+comments.
3.2.3 Applying Heuristics
In the initial phase of our analysis, we experimented with various sizes of n-gram to determine which configurations offered the most valuable insights for detecting recurring patterns. In contrast to the methodology employed by  Tufano et al. [9], which used 2- or 3-grams, we found that a maximum size of 5-n-gram provided sufficient context or relevant information to accurately capture the role of GPT in GPT-assisted pull requests. We then proceeded to automatically extract both forward and backward n-grams, up to a maximum of 5-grams [32], containing the term GPT, while simultaneously applying a set of heuristics designed to reduce the occurrence of false positives. For instance, when analyzing the PR body, we encountered the sentence: ”Instead of incorporating muP into GPT-NeoX we are going to move these changes to [our fork of their repo]” [^1] https://github.com/EleutherAI/gpt-neox/pull/1061#issuecomment-1771635901. From this, we extracted the backward n-gram ”of incorporating muP into GPT-NeoX” and the forward n-gram ”GPT-NeoX we are going to”; the application of heuristic H4 identified GPT-NeoX as a reference to the PR name, thereby classifying this PR as non-GPT-assisted. Additionally, we applied heuristics to filter and automatically eliminate false positives, including cases where GPT is mentioned in a PR but not used for assistance, as well as instances where a PR may appear GPT-assisted but is actually unrelated (e.g., mentions of GPT in the context of non-code-related discussions, such as research or tool integration). These measures ensured that only relevant PRs, where GPT is actively employed for code review assistance, were included in the analysis.
3.2.4 Inspection
To ensure the reliability of our heuristics, two authors independently examined 200 randomly chosen PRs by carefully reviewing the labels generated during the execution of the heuristics. They then discussed unambiguous cases where the labeling was unclear to reach an agreement, such as whether the PR used GPT as an assistant. In the final step, the heuristics were adjusted and reapplied to improve accuracy and consistency. We performed three rounds of refinement/reapply until we obtained an accuracy of 95% as an acceptable agreement score in the labeling of PRs.
Table 1: GPT PRs During a Round of Sampling Analysis
Filter Assistance Non Assistance
Project name 4 21
PR title 6 19
Modified files 23 58
PR body 1 24
Total 34 122
3.2.5 Combining the Heuristics Outputs
We reapplied the heuristics to the entire dataset to discard a higher number of false positives. All PRs were labelled GPT-assisted or GPT-non-assisted after applying the heuristics (as detailed in Table 2). Subsequently, one of the authors examined the dataset for true positive cases by manually reviewing the final labels. Ultimately, we split the dataset into two separate groups: “GPT-assisted” and “GPT-non-assistance” PRs. The table 3 displays the distribution of GPT-assisted and GPT-non-assisted PR in the dataset.
@@ 3.2.5 Combining the Heuristics Outputs
Table 3: Final Dataset: Assistance and Non-Assistance GPT PR Status Summary
Status Assistance Non Assistance
Open 54 38
Closed 179 219
Merged 1367 1036
Total 1600 1293
3.3 Mapping GPT-assisted to Corresponding PR Phases
In this automatic step, our goal is to define the different phases of the code review process as previously performed by Kudrjavets et al. [5]. We report descriptive statistics to quantify the extent of GPT usage in each phase, depending on whether non-productive phases are present in the review workflow.

### v2

@@ Abstract
Large language models (LLMs) have recently gained prominence in the field of software development, significantly boosting productivity and simplifying teamwork. Although prior studies have examined task-specific applications, the phase-specific effects of LLM assistance on the efficiency of code review processes remain underexplored. This research investigates the effect of GPT on GitHub pull request (PR) workflows, with a focus on reducing resolution time, optimizing phase-specific performance, and assisting developers. We curated a dataset of 25,473 PRs from 9,254 GitHub projects and identified GPT-assisted PRs using a semi-automated heuristic approach that combines keyword-based detection, regular expression filtering, and manual verification until achieving 95% labeling accuracy. We then applied statistical modeling, including multiple linear regression and Mann–Whitney U test to evaluate differences between GPT-assisted and non-assisted PRs, both at the overall resolution level and across distinct review phases. Our research has revealed that early adoption of GPT as assistant can boost the effectiveness of the PR process, leading to considerable time savings at various stages. Our findings suggest that GPT-assisted PRs reduced median resolution time by more than 61%. We discovered that utilizing GPT can reduce the review time by 66.7%, and the waiting time before acceptance by 87.5%. Analyzing a sample of 300 GPT-assisted PRs, we discovered that developers predominantly use GPT for code optimization (60%), bug fixing (26%), and documentation updates (12%). Our findings highlight the potential of GPT to accelerate code reviews, improve code quality, and address persistent bottlenecks. This research sheds light on the impact of the GPT model on the code review process, offering actionable insights for software teams seeking to enhance workflows and promote seamless collaboration.
@@ 1 Introduction
Modern software development is heavily based on collaborative workflows, with Pull Requests (PR) serving as a central mechanism to integrate code changes [gousios2014exploratory, sadowski2018modern, bacchelli2013expectations]. A PR goes through multiple phases, including submission, review, change, and merging, each requiring coordination between contributors and reviewers [gousios2014exploratory]. Effective PR management is crucial to maintaining software quality, but it remains a bottleneck in many projects due to long review times, inconsistent feedback, and inefficient communication [zhang2022pull]. Developers often experience delays due to unclear review comments, excessive revisions, and prolonged waiting times before changes are accepted [kudrjavets2022mining].
Inefficiencies in the software development cycle, such as long review times, can lead to frustration among team members, decreased productivity, and even project delays [baysal2016investigating]. As codebases become increasingly intricate, there is a growing need for sophisticated automation tools to facilitate the code review process and improve collaboration and teamwork [zheng2024code]. In parallel, LLMs, such as Generative Pre-trained Transformers (GPTs) [vaswani2017attention] have rapidly evolved, demonstrating strong capabilities in code generation, bug detection, and documentation assistance [tufano2024unveiling, ebert2023generative]. These models are capable of suggesting enhancements to existing code, resolving unclear logical structures, and automating repetitive tasks, thereby becoming indispensable tools for software engineers.
Despite the increasing adoption of LLM-powered tools in development workflows, a systematic investigation into their impact on PR processes is lacking. Previous research on the use of LLM in the PR context [tufano2024unveiling, grewal2024analyzing, chouchen2024so, rasheed2024ai] has focused mainly on task-specific applications, such as commit message generation [xiao2024generative] and documentation support [davila2024tales], while neglecting to explore their phase-specific impact across the entire PR lifecycle. Furthermore, while LLM-generated suggestions can be beneficial, it is necessary to quantify their effectiveness in improving review speed and reducing PR merge time.
@@
Toward this goal, we examine the PRs that software developers actively engage with GPT in conversational exchanges to seek recommendations, clarifications, or solutions related to code review tasks. We refer to these as GPT-assisted PRs throughout the paper. For example, in this GPT-assisted PR [pullrequest2059], approximately 99% of the code was generated by GPT-4, as the developer requested ChatGPT to enhance a Ruby plugin that enables Jekyll to fetch posts from an external RSS feed when rss_url is specified. In contrast, GPT non-assisted PRs are those where developers either do not use GPT at all or use it strictly for implementation tasks (i.e., making an API call to a GPT model), without relying on it for code review insights or decision making. Our research investigates how GPT-assisted PRs influence the duration and quality of the code review process. For this purpose, we start by curating a dataset of 25,473 PRs from 9,254 projects on GitHub. Next, we extract six features (project_age, no_commits, files_added, files_deleted, files_changed, time_to_merge) to characterize projects, PRs, and code reviews. We then analyze how these factors influence the review and resolution times [moreira2021factors]. Finally, we also investigate which development tasks involve the use of GPT-assisted PRs throughout the PR phases, aiming to identify where such assistance is most effectively integrated. In summary, our aim is to answer the following three research questions:
@@ 2 Related Work
Challenges in Modern Code Reviews (MCR) and Collaboration. Code reviews are a critical aspect of modern software development, ensuring code quality, maintainability, and adherence to best practices. However, the process comes with several challenges that can affect efficiency, collaboration, and overall development workflow. For example, long waiting times where PRs often remain unreviewed for extended periods, slowing down development. kudrjavets2022mining reported that between 29% and 63% of the overall code review lifetime is spent waiting for the accepted changes to merge. bacchelli2013expectations reveal that scheduling and time issues also appeared challenging; yet, understanding the code and comments takes most of the review time. According to these authors, code reviews tend to be less focused on detecting defects than expected and, instead, provide additional advantages, such as knowledge transfer, increased team awareness, and the creation of alternative solutions to problems. davila2021systematic also reported that reviewers have found it difficult to understand the code, which prompts the need for enhancements to visualize code changes to facilitate understanding. However, a recent study has shown that ChatGPT is an alternative to generate some kind of explanations during MCR [widyasari2023explaining]. davila2024tales identified five challenges to adopting generative AI in MCR: trustworthiness, lack of context, misleading review comments, security and privacy, and token limits. Although prior studies have explored aspects of the use of GPT in software development, they do not provide a systematic analysis of its impact on waiting times and efficiency in all phases of the PR workflow. Our study addresses this gap by focusing on its influence during specific code review phases.
@@
Factors Affecting PR Lifetime and Latency. Studying the factors that influence PR latency is crucial for improving software development efficiency, collaboration, and overall code quality. Understanding what causes delays in PR review and merging can lead to optimizations that benefit both developers and organizations [zhang2022pull], [rose2017towards], [moreira2021factors], [khatoonabadi2024predicting], and [khatoonabadi2023wasted]. zhang2022pull identified that the number of commits at open time is negatively correlated with PR merging but shifts to a positive correlation at close time. rose2017towards built a Multiple Linear Regression Model (MLR) that indicates that the PR size has a statistically significant effect on the review time, as the positive value of the regression coefficient implies that as the PR size increases, the review time for PRs also increases. moreira2021factors further emphasized that PRs possess some structural characteristics, such as the number of modified files, the number of commits, and the number of changed lines of code. moreira2021factors concluded that the combination of structural characteristics may increase the chances of PRs having a very short and long lifetime.
@@
Although prior research explores factors affecting PR latency (e.g., size, number of commits), there is limited research into how GPT can have an effect on PR resolution time. khatoonabadi2024predicting developed a machine learning model to predict the latency of the initial responses of the maintainers and contributors after receiving the first feedback from the maintainer. These authors highlighted that the timing of submissions and feedback are factors that have a significant influence on response waiting times. Their approach, which provides estimated waiting times, helps open-source projects collaborate better by allowing maintainers and contributors to anticipate and address potential delays in PR reviews. This study complements the aforementioned work by offering empirical insights into GPT usage beyond commit messages, specifically examining detailed time reductions in key phases such as review and change.
@@
The Application of LLMs in Code Reviews. LLM-powered tools can offer services designed to streamline various code review-related development tasks. Research has been conducted on the impact of AI-driven generative tools in this context [tufano2024unveiling], [xiao2024generative], [davila2024tales], [rasheed2024ai], [vijayvergiya2024ai], [yu2024fine], [hao2024empirical], [das2024investigating]. For instance, davila2024tales conducted a literature review on the application of AI-based generative tools in a code review context and highlighted that AI-based tools can positively impact code review efficiency by reducing reviewer workload and review duration and increasing the likelihood of acceptance in GitHub projects. xiao2024generative observed that PRs assisted by Copilot have a 1.57 times higher likelihood of being merged compared to those not assisted by Copilot for PR. tufano2024unveiling reported how developers use ChatGPT to automatically generate a commit message and a PR/issue description. rasheed2024ai highlighted the potential of LLM-powered AI agents to considerably improve conventional code review procedures. These agents provide a twofold advantage: they enhance code quality and improve developer learning.
@@
vijayvergiya2024ai shared the experience in developing, implementing, and evaluating AutoCommenter, a LLM-powered code review tool that enforces coding best practices and has a positive impact on the developer’s workflow. yu2024fine proposed Carllm (Comprehensibility of Automated Code Review using LLMs) that improves the comprehensibility of automated code review. hao2024empirical studied the role of ChatGPT in collaborative coding by analyzing the shared conversation between developers and ChatGPT in GitHub PRs and issues, finding that the most frequent requests involve code generation, conceptual understanding, how-to guidance, issue resolution, and code review. Studies on GPT have examined particular tasks (e.g., code generation, commit message generation, or bug fixing), but lack a holistic view of its phase-specific impacts during the PR life cycle.
cihan2025automated conduct an empirical study on the use of LLM-based automated code review in the industry. They analyze 4,335 PRs across three projects. Although approximately 74% of automated comments were applied, the tool slightly improved code quality, increased bug detection, and promoted best practices, but also led to longer PR closure times and occasional irrelevant suggestions. The study shows that automated reviews can help developers, but they do not replace human judgment. Effectiveness depends on the context of the project and current workflows. ramesh2025automated present a case study on an LLM-augmented code review tool used by Ericsson. The tool combines static program analysis with prompt engineering to deliver context-aware reviews in a live workflow. Their evaluation suggests that integrated LLM-based reviews can reduce manual effort and support developer workflows. However, effective prompting, tool integration, and risk mitigation remain key challenges. Previous research has explored the use of large language models for automated code review in software development, but human involvement is still essential to ensure contextual understanding. This study aims to address this limitation by investigating the impact of GPT on specific phases of the PR workflow, specifically the review and the waiting phase. It is important to note that this study does not propose a new approach for automatically generating code reviews using LLMs.
@@ 3 Study Setup
Figure 1: An overview of our study setup that describes a semi-automated approach for identifying, detecting, and mapping GPT-assisted PRs to the phases in which GPT is used.
In this paper, we investigate how incorporating GPT as an assistance tool influences the code review process. In particular, it focuses on scenarios where developers interact with GPT to obtain feedback or suggestions during PR reviews, while excluding PRs generated automatically by GPT. Specifically, we examine how GPT influences the overall PR resolution time and how its effects vary between different phases of the review workflow. We also aim to identify the types of review tasks for which developers actively seek GPT support. Figure 1 presents an overview of our study design. We begin with an automated procedure to identify potential GPT-assisted PRs, followed by a semi-automated verification step to confirm the actual use of GPT during the review. Finally, we map the verified GPT-assisted PRs to their corresponding review phases. The remainder of this section provides a detailed explanation of each step in our methodology.
@@ 3.1 Searching Potential GPT-assisted PRs
To identify relevant PRs for our analysis, we rely on GitHub, which adopts a PR-based approach to submit, review, and merge code contributions [davila2021systematic]. We use the GitHub REST API [Github-API] to retrieve a broad set of candidate PRs in which the assistant of a GPT model could have been used during the review process.
@@ 3.1.1 Searching
Following a similar strategy to that of  tufano2024unveiling, we conduct keyword-based searches for mentions of ”GPT” in four main PR fields, obtaining a total of 81,234 PRs.
@@ 1. $\blacksquare$
    Project Name: Although some GPT-focused projects may not actually use GPT for PR reviews, we still record these mentions before filtering out irrelevant cases [pullrequest43].
@@ 2. $\blacksquare$
    PR Title: The use of GPT may be evident in the PR title (e.g., refactored using GPT suggestions [pullrequest64]).
@@ 3. $\blacksquare$
    Body+Comments: A PR description and/or comments can refer to GPT explicitly (e.g., code generated by GPT [pullrequest16]).
@@ 4. $\blacksquare$
    File(s) Name: References in filenames (e.g., chatgpt_review.md [pullrequest1]) can indicate GPT-generated artifacts.
@@ 3.1.2 Filtering
To avoid projects with minimal activity or that lack meaningful collaboration, we keep only projects that meet the following criteria: (1)  repositories that have at least 10 stars [borges2018s], to remove toys or trivial projects. (2) at least two contributors are involved in the review process, to exclude self-reviewed PRs. (3) PRs that contain more than one event in the lifetime.
@@ 3.2 Identifying GPT-assisted PRs
In this semi-automated step, our goal is to define the heuristics that indicate the use of GPT to classify PRs as GPT-assisted or GPT-non-assisted. By doing this, we exclude false positives (i.e., PRs that mentioned GPT as discussed in Section 3.1 but did not use GPT to assist in the code review process). This step ensures that the PRs used for our analysis are PRs where GPT models are used in the code review process, rather than just being part of the project code.
@@ 3.2.1 Defining Heuristics
To enable the identification of patterns related to GPT usage in code review processes, we randomly selected 450 GitHub PRs from the initial dataset with a confidence level of 95% and a margin error of 5%. We ensure an even distribution among the three contributing authors to examine indicators of GPT assistance during review activities.
Three authors independently inspected 150 unique PRs in three separate rounds, analyzing project names, PR titles, modified files, and PR bodies. In Round 1, each annotator reviewed the first set of 150 PRs and cross-validated our observations to ensure that the identified patterns consistently reflected GPT-assisted contributions. In particular, most PRs were ultimately classified as GPT-non-assisted, as GPT was often integrated into the software (e.g., through GPT APIs) but not explicitly used to support the code review itself. Table 5 in the appendix presents the findings of Round 1.
In Rounds 2 and 3, the annotators applied the same procedure to two additional sets of 150 PRs each. After each round, the reviewers met to consolidate recurring patterns, clarify ambiguous cases, and resolve classification disagreements through discussion until consensus was reached. For relevance labeling, disagreements occurred in 4% of the cases. Krippendorff’s $\alpha=0.77$ [krippendorff2011] indicates substantial inter-annotator reliability. This suggests that the evaluators were largely consistent in their judgments, providing confidence in the quality of the relevance annotations.
We defined four heuristics, as outlined in Table 1, to analyze and label the presence of GPT-assisted PRs. These heuristics are reinforced by a regular expression-based technique which automatically classifies the collected PR instances as either GPT-assisted or GPT-non-assisted based on the identified patterns: H1: Project name, H2: PR title, H3: Modified files, H4: PR body+comments.
@@
3.2.2 Label PRs using Heuristics
Initially, we experimented with various sizes of n-gram to determine which configurations offered the most valuable insights for detecting recurring patterns. In contrast to the methodology employed by  tufano2024unveiling, which used 2- or 3-grams, we found that a maximum size of 5-n-gram provided sufficient context or relevant information to accurately capture the role of GPT in GPT-assisted pull requests. We then proceeded to automatically extract both forward and backward n-grams, up to a maximum of 5-grams [jurafsky2023slp3], containing the term GPT, while simultaneously applying a set of heuristics designed to reduce the occurrence of false positives. For instance, when analyzing the PR body, we encountered the sentence: ”Instead of incorporating muP into GPT-NeoX we are going to move these changes to [our fork of their repo]” [^1] https://github.com/EleutherAI/gpt-neox/pull/1061#issuecomment-1771635901. From this, we extracted the backward n-gram ”of incorporating muP into GPT-NeoX” and the forward n-gram ”GPT-NeoX we are going to”; the application of heuristic H4 identified GPT-NeoX as a reference to the PR name, thereby classifying this PR as GPT non-assisted. Additionally, we applied heuristics on the sample of 450 PRs (defined in section 3.2.1) to filter and automatically eliminate false positives, including cases where GPT is mentioned in a PR but not used for assistance, as well as instances where a PR may appear GPT-assisted but is actually unrelated (e.g., mentions of GPT in the context of non-code-related discussions, such as research or tool integration). These measures ensured that only relevant PRs, where GPT is actively used for code review assistance, were included in the analysis.
To ensure the reliability of our heuristics, two authors independently examined the labeling on the 450 PRs by carefully reviewing the labels generated during the execution of the heuristics. In the final step, the heuristics were adjusted and reapplied to improve accuracy and consistency. We performed three rounds of refinement/reapply until we obtained an accuracy of 95% as an acceptable agreement score in the labeling of PRs. Table 6 in the appendix presents the increase in accuracy as more heuristics were refined.
3.2.3 Applying the Heuristics
We applied the heuristics to the entire dataset of 25,473 PRs to label them as either GPT-assisted or non-assisted. Subsequently, one of the authors examined the dataset for true positive cases by manually reviewing the final labels. Ultimately, we split the dataset into two separate groups: GPT-assisted and GPT-non-assistance PRs. Of the 1,600 GPT-assisted PRs, 54 remain open, 179 have been closed without merging, and 1,367 have been successfully merged. This distribution highlights that the majority of GPT-assisted PRs are ultimately integrated into the codebase, reflecting the potential influence of GPT on the completion and acceptance of code contributions. The table 7 displays the distribution of GPT-assisted PRs in the dataset.
3.3 Mapping Events to PR Phases
In this automatic step, our goal is to define the different phases of the code review process as previously performed by kudrjavets2022mining. We report descriptive statistics to quantify the extent of GPT usage in each phase, depending on whether non-productive phases are present in the review workflow.
3.3.1 Finding Similarity GPT-non-assisted PRs with GPT-Assisted PRs
To have a fair comparison between the GPT-assisted and GPT-non-assisted PRs, we quantify the similarity between them applying the Manhattan distance [malkauthekar2013analysis, ye2021recommending] on the same factors as applied in RQ1 4.1. To mitigate the presence of outliers or bias in the factor, we rescale no_commits, PR_size, no_changed_files, project_age to a range within [0, 1], applying the max-min method [ye2021recommending]. The two well-known metrics (the Manhattan and Euclidean distance produce very similar results (Spearman correlation $\rho=0.98$ [JSSv094i12] when computing the distances between factors), we finally test the differences between the PRs supported by or without GPT by getting the minimum distance value from the Manhattan distance.